OpenStudy (anonymous):

HELP!!Solve 1 2 - + - 5 5 3 --- 5 ------------------------------------------ 2 --- 5 ----------------------------------------- 1 --- 5 ----------------------------------------- 3 --- 10

OpenStudy (anonymous):

really

OpenStudy (anonymous):

thats so easy r u a elementry or kindergarden student

OpenStudy (anonymous):

Its a learning site not a hating site smart one

OpenStudy (anonymous):

\[\frac{1}{5}+\frac{2}{5} = \frac{1+2}{5}\]

OpenStudy (anonymous):

so 3/5

OpenStudy (anonymous):

yes that is correct.

OpenStudy (anonymous):

@Mommasboy nevermind